The product of two consecutive integers is 5 less than the square of the lesser integer. Find the greater of the two integers.
Found 2 solutions by richwmiller, ikleyn:
Answer by richwmiller(17219)

Oops! I read 2 consecutive even integers!

Answer by ikleyn(52835)

.
x*(x+1) = x^2 - 5


x^2 + x = x^2 - 5


x = -5  is the lesser number.


x = -4  is your ANSWER.


CHECK.  a)  (-5)*(-4) = 20.    b)  (-5)*(-5) - 5 = 25 - 5 = 20.

        ! Correct !

Solved, answered and completed. And explained.